A solution containing 6 g of benzoic acid is 80 g of ethere has a vapour pressure of 40mmHg at 293 k given that the vapour pressure of ethereal same temperature is 442mmHg calculate the molar mass of benzoic acid 
 

Grade:12

1 Answers

Utkarsh mishra
5 years ago
According to raoult's law for dilute solution
 = P - Ps /P°= moles of solute / moles of solvent 
=}442-410/442 = 6/M/50/70 
}0.072= 6/M/5/7
M= 122.65 molecular mass